OpenStudy (anonymous):

Okay, so if i have x^2+4x-72=5x, i can subtract the 5x to get x^2-1x-72=0, right? How do you find the other numbers to go in the ( ) with the x?

OpenStudy (phi):

Here's what I do. Look at the third number -72

OpenStudy (anonymous):

Okay...

OpenStudy (phi):

now the hard part, find its factors 1 x 72 2 x 36

OpenStudy (phi):

3*26 4*18

OpenStudy (phi):

6*12 8*9

OpenStudy (phi):

Now because we have negative 72,(-72) find a pair when you subtract them, give you the second number in your equation ( -1 in front of the x)

OpenStudy (anonymous):

as before, the two numbers are 8 and -9. So (x+8)(x-9), you get x=-8, x=9.

OpenStudy (anonymous):

Oh! So you just find what factors of the last number give you the second number, then you switch the signs?

OpenStudy (phi):

As you notice 8-9= -1 so those are our numbers so the factors will be (x+8)(x-9)

OpenStudy (phi):

and if you are not sure. double check the answer by multiply it out (x+8)(x-9) = x^2 +8x -9x -72 = x^2 -x -72

OpenStudy (phi):

Now if the 3rd number were positive, we would have to find the sum of a pair for example, x^2 + 17x +72 we would use 8 and 9, because they add to 17. The factors would be (x+8)(x+9)

OpenStudy (anonymous):

Oh, so the 3rd number shows you wether you will have negatives in your factors or not?

OpenStudy (phi):

Yes!

OpenStudy (anonymous):

not to butt in but in enlgish what you are looking for is two numbers whose produce is -72 and whose sum is -1 i.e. when you multiply you get -72 and when you add you get -1 just thought i would mention it

OpenStudy (anonymous):

Do you always have to subtract to find get your second number out of the factors, or does that also depend on some other number?

OpenStudy (phi):

If the third number is negative, you must subtract the pair to get the 2nd number.

OpenStudy (anonymous):

And if its positive you add?

OpenStudy (phi):

example x^2+20x - 21 factors 1 * 21 3*7 21 - 1 = 20 so that is what we want and the answer is (x+21)(x-1)

OpenStudy (phi):

and if the 3rd number is positive you add the pair. example x^2 +22x +21 we use the pair 1,21 because 1+21= 22 (and the 3rd number 21 is positive) (x+1)(x+21)

OpenStudy (anonymous):

What if the second number is negative and the third number is positive?

OpenStudy (phi):

you end up with complex numbers, and have to use the quadratic formula. You may not have learned of it?

OpenStudy (anonymous):

Is the quadratic formula the one thats really long and confusing?

OpenStudy (phi):

not for me!

OpenStudy (phi):

\[a ^{2}+bx+c\] the roots are \[x= \frac{-b \pm \sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (anonymous):

... That looks pretty long and confusing...

OpenStudy (phi):

that should be ax^2

OpenStudy (phi):

practice helps

OpenStudy (anonymous):

Haha, alright, ill practice late on those types of problems. (: Thanks for your help! I actually understand it.

OpenStudy (phi):

Yep, I'd leave it for another day... no need to overload, and fry your brains!

OpenStudy (anonymous):

Haha, i wouldnt want that to happen! (:

OpenStudy (phi):

Hi, I have to correct myself. You asked What if the second number is negative and the third number is positive? What you do is still "add" but use negative numbers. For example x^2 -4x +3 the factors for 3 is just the one pair 1*3, add to get 4. But the 2nd number is negative, so we should really add -1 - 3 = -4. So our two numbers are -1 and -3, and our factors are (x-1)(x-3) which multiples out to x^2 -x -3x +3 = x^2 -4x +3 Hope that's clear....